# HIGHT: Hierarchical Graph Tokenization for Graph-Language Alignment

### v1

@@ B.1 Details of the PubChem Dataset
B.2 Details of HiPubChem Dataset
@@
B.3 Details of Property Prediction Dataset
The task of molecular property prediction mainly aims to predict certain biochemical or physical properties of molecules. Usually, these properties have a close relation with the molecular substructures (i.e., functional groups) [Bohacek et al., 1996]. In this work, we consider the scenarios of both binary classification based and the regression based molecular property prediction, and the datasets are mainly derived from MoleculeNet [Wu et al., 2017].
@@ B.3 Details of Property Prediction Dataset
For regression, we adopt the instruction tuning data from Mol-Instructions [Fang et al., 2024]. The regression based property prediction focuses on predicting the quantum mechanics properties of the molecules. The 1D sequence information in this task is given by SELFIES [Krenn et al., 2019]. The original data is sourced from the QM9 subset of the MolculeNet [Wu et al., 2017]. There are three subtasks: (i) Highest occupied molecular orbital (HOMO) energy prediction; (ii) Lowest occupied molecular orbital (LUMO) energy prediction; (iii) and HUMO-LUMO gap energy prediction. Some examples of the regression based property prediction dataset are given in Table 10.
B.4 Details of Reaction Prediction Dataset
We adopt three chemical reaction related tasks from Mol-Instructions [Fang et al., 2024]: Forward reaction prediction, reagent prediction, and retrosynthesis prediction. The input and output contain 1D sequence information given by SELFIES [Krenn et al., 2019]. Some examples of the Mol-Instructions datasets are given in Table 11, where the SELFIES represented molecules are denoted as “<SELFIES>” for clarity.
@@ B.4 Details of Reaction Prediction Dataset
Forward Reaction Prediction Question: With the provided reactants and reagents, propose a potential product.<SELFIES>
Answer: <SELFIES>
Reagent Prediction Question: Please suggest some possible reagents that could have been used in the following chemical reaction. The reaction is <SELFIES>
Answer: <SELFIES>
Retrosynthesis Prediction Question: Please suggest potential reactants for the given product. The product is: <SELFIES>
Answer: <SELFIES>
@@
B.5 Details of Molecular Description Dataset
For the molecular description task, we adopt a widely used dataset ChEBI-20 [Edwards et al., 2021]. Based on the molecules from PubChem, Edwards et al. [2021] collected the Chemical Entities of Biological Interest (ChEBI) [Hastings et al., 2015] annotations of the molecules, which are the descriptions of molecules. We transform the task into the instructions, and present some samples in Table 12. The authors collect $33,010$ molecule-text pairs and split them into training ($80$%), validation ($10$%), and testing ($10$%) subsets. We mainly adopt the original training split to tune the model and evaluate the tuned model on the original test split.
@@ B.5 Details of Molecular Description Dataset
B.6 Details of MotifHallu Dataset
@@
Then, we leverage RDKit [Landrum, 2016] to detect the existence of the left $38$ valid functional groups within a molecule. We consider $3,300$ molecules from ChEBI-20 test split [Edwards et al., 2021], and adopt the query style as for large vision-language models [Li et al., 2023c] that queries the existence of specific functional group one by one:
@@
During the evaluation, we detect whether the LGLM gives outputs meaning “Yes” or “No” following the practice in [Li et al., 2023c]. For each molecule, we construct questions with positive answers for all kinds of functional groups detected in the molecule, and questions with negative answers for randomly sampled $6$ functional groups from the $38$ common functional groups in RDKit. The construction finally yields $23,924$ query answer pairs about the existence of functional groups in the molecule. While it is easy to scale up MotifHallu by automatically considering more molecules and a broader scope of functional groups, we find that the current scale is already sufficient to demonstrate the hallucination phenomena in LGLMs.
@@
Appendix C Details of Experiments
@@
We implement the GNN tokenizer/encoder based on the same GNN backbone, which is a $5$-layer GIN [Xu et al., 2019]. The hidden dimension is $300$. For the node-centric tokenization, we employ the VQVAE GNN tokenizer from Mole-BERT [Xia et al., 2023] and adopt self-supervised learning tasks from the official Mole-BERT implementation. [^7] https://github.com/junxia97/Mole-BERT For HIGHT, we train the VQVAE with the self-supervised learning tasks from [Zang et al., 2023] based on the official implementation. [^8] https://github.com/ZangXuan/HiMol Meanwhile, we set the hyperparameters of GNN tokenizer training the same as those recommended by [Xia et al., 2023, Zang et al., 2023].
@@
For the LoRA adapters, we use a LoRA rank of $64$ and a scaling value $\alpha$ of $16$ [Hu et al., 2022] for all methods and tasks.
For the base LLM, we mainly adopt vicuna-v-1.3-7B [Chiang et al., 2023]. The overall scale of parameters is around $6.9$B.
@@
For the evaluation of classification-based property prediction, we adopt the ROC-AUC following the common practice [Wu et al., 2017].
For the evaluation of regression-based property prediction, we adopt the Mean Absolute Error (MAE) following the common practice [Fang et al., 2024].
For the evaluation of molecular description, we adopt BLEU-2, BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, and METEOR following the common practice [Papineni et al., 2002, Lin, 2004, Edwards et al., 2021]. To improve the reliability of the evaluation, the metrics are computed based on the tokenizer scibert_scivocab_uncased of SciBERT [Beltagy et al., 2019].
We follow the common practice to evaluate models for the tasks of chemical reaction predictions [Fang et al., 2024]. We adopt linguistic metrics such as BLEU [Papineni et al., 2002], ROUGE-L [Lin, 2004], METEOR [Banerjee and Lavie, 2005] and Levenshtein scores [Yujian and Bo, 2007]. Meanwhile, we also validate the validity of the generated molecular sequences with RDKit [Landrum, 2016]. In addition, several molecular similarity measures are also leveraged. Specifically, we present the MAE of the RDKit, MACCS, and Morgan fingerprints to assess the semantic similarity of the generated compounds and the ground truth ones [Durant et al., 2002, Schneider et al., 2015].
As for the MotifHallu, in order to avoid the drawbacks that LGLMs may output answers that do not follow the instructions, we compare the loss values by feeding the answers of “Yes” and “No”, and take the one with a lower autoregressive language modeling loss as the answer. Following the practice in LVLMs, we present the F1 scores, accuracies, and the ratio that the model answers “Yes” [Li et al., 2023c]. Given the severe imbalance of positive and negative samples, we separately report the F1 scores for positive and negative classes.
@@
We implement our methods with PyTorch 11.3 [Paszke et al., 2019]. We run experiments on Linux Servers with NVIDIA V100 and NVIDIA A100 (40G) graphics cards with CUDA 11.7.

### v2

@@ C.1 Details of the PubChem Dataset
C.2 Details of HiPubChem Dataset
@@
C.3 Details of Property Prediction Dataset
The task of molecular property prediction mainly aims to predict certain biochemical or physical properties of molecules. Usually, these properties have a close relation with the molecular substructures (i.e., functional groups) (Bohacek et al., 1996). In this work, we consider the scenarios of both binary classification based and the regression based molecular property prediction, and the datasets are mainly derived from MoleculeNet (Wu et al., 2017).
@@ C.3 Details of Property Prediction Dataset
For regression, we adopt the instruction tuning data from Mol-Instructions (Fang et al., 2024). The regression based property prediction focuses on predicting the quantum mechanics properties of the molecules. The 1D sequence information in this task is given by SELFIES (Krenn et al., 2019). The original data is sourced from the QM9 subset of the MolculeNet (Wu et al., 2017). There are three subtasks: (i) Highest occupied molecular orbital (HOMO) energy prediction; (ii) Lowest occupied molecular orbital (LUMO) energy prediction; (iii) and HUMO-LUMO gap energy prediction. Some examples of the regression based property prediction dataset are given in Table 13.
C.4 Details of Reaction Prediction Dataset
We adopt three chemical reaction related tasks from Mol-Instructions (Fang et al., 2024): Forward reaction prediction, reagent prediction, and retrosynthesis prediction. The input and output contain 1D sequence information given by SELFIES (Krenn et al., 2019). Some examples of the Mol-Instructions datasets are given in Table 14, where the SELFIES represented molecules are denoted as “¡SELFIES¿” for clarity.
@@ C.4 Details of Reaction Prediction Dataset
Forward Reaction Prediction Question: With the provided reactants and reagents, propose a potential product.¡SELFIES¿
Answer: ¡SELFIES¿
Reagent Prediction Question: Please suggest some possible reagents that could have been used in the following chemical reaction. The reaction is ¡SELFIES¿
Answer: ¡SELFIES¿
Retrosynthesis Prediction Question: Please suggest potential reactants for the given product. The product is: ¡SELFIES¿
Answer: ¡SELFIES¿
@@
C.5 Details of Molecular Description Dataset
For the molecular description task, we adopt a widely used dataset ChEBI-20 (Edwards et al., 2021). Based on the molecules from PubChem, Edwards et al. (2021) collected the Chemical Entities of Biological Interest (ChEBI) (Hastings et al., 2015) annotations of the molecules, which are the descriptions of molecules. We transform the task into the instructions, and present some samples in Table 15. The authors collect $33,010$ molecule-text pairs and split them into training ($80$%), validation ($10$%), and testing ($10$%) subsets. We mainly adopt the original training split to tune the model and evaluate the tuned model on the original test split.
@@ C.5 Details of Molecular Description Dataset
C.6 Details of MotifHallu Dataset
@@
Then, we leverage RDKit (Landrum, 2016) to detect the existence of the left $38$ valid functional groups within a molecule. We consider $3,300$ molecules from ChEBI-20 test split (Edwards et al., 2021), and adopt the query style as for large vision-language models (Li et al., 2023c) that queries the existence of specific functional group one by one:
@@
During the evaluation, we detect whether the LGLM gives outputs meaning “Yes” or “No” following the practice in (Li et al., 2023c). For each molecule, we construct questions with positive answers for all kinds of functional groups detected in the molecule, and questions with negative answers for randomly sampled $6$ functional groups from the $38$ common functional groups in RDKit. The construction finally yields $23,924$ query answer pairs about the existence of functional groups in the molecule. While it is easy to scale up MotifHallu by automatically considering more molecules and a broader scope of functional groups, we find that the current scale is already sufficient to demonstrate the hallucination phenomena in LGLMs.
@@
Appendix D Details of Experiments
@@
We implement the GNN tokenizer/encoder based on the same GNN backbone, which is a $5$-layer GIN (Xu et al., 2019). The hidden dimension is $300$. For the node-centric tokenization, we employ the VQVAE GNN tokenizer from Mole-BERT (Xia et al., 2023) and adopt self-supervised learning tasks from the official Mole-BERT implementation. [^8] https://github.com/junxia97/Mole-BERT For HIGHT, we train the VQVAE with the self-supervised learning tasks from (Zang et al., 2023) based on the official implementation. [^9] https://github.com/ZangXuan/HiMol Meanwhile, we set the hyperparameters of GNN tokenizer training the same as those recommended by (Xia et al., 2023; Zang et al., 2023).
@@
For the LoRA adapters, we use a LoRA rank of $128$ and a scaling value $\alpha$ of $256$ for molecular property prediction (classification) in order to better fit with the task, and use a LoRA rank of $64$ and a scaling value $\alpha$ of $16$ for all the remaining methods and tasks.
For the base LLM, we mainly adopt vicuna-v-1.3-7B (Chiang et al., 2023). The overall scale of parameters is around $6.9$B.
@@
For the evaluation of classification-based property prediction, we adopt the ROC-AUC following the common practice (Wu et al., 2017).
For the evaluation of regression-based property prediction, we adopt the Mean Absolute Error (MAE) following the common practice (Fang et al., 2024).
For the evaluation of molecular description, we adopt BLEU-2, BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, and METEOR following the common practice (Papineni et al., 2002; Lin, 2004; Edwards et al., 2021). To improve the reliability of the evaluation, the metrics are computed based on the tokenizer scibert_scivocab_uncased of SciBERT (Beltagy et al., 2019).
We follow the common practice to evaluate models for the tasks of chemical reaction predictions (Fang et al., 2024). We adopt linguistic metrics such as BLEU (Papineni et al., 2002), ROUGE-L (Lin, 2004), METEOR (Banerjee & Lavie, 2005) and Levenshtein scores (Yujian & Bo, 2007). Meanwhile, we also validate the validity of the generated molecular sequences with RDKit (Landrum, 2016). In addition, several molecular similarity measures are also leveraged. Specifically, we present the MAE of the RDKit, MACCS, and Morgan fingerprints to assess the semantic similarity of the generated compounds and the ground truth ones (Durant et al., 2002; Schneider et al., 2015).
As for the MotifHallu, in order to avoid the drawbacks that LGLMs may output answers that do not follow the instructions, we compare the loss values by feeding the answers of “Yes” and “No”, and take the one with a lower autoregressive language modeling loss as the answer. Following the practice in LVLMs, we present the F1 scores, accuracies, and the ratio that the model answers “Yes” (Li et al., 2023c). Given the severe imbalance of positive and negative samples, we separately report the F1 scores for positive and negative classes.
@@
We implement our methods with PyTorch 11.3 (Paszke et al., 2019). We run experiments on Linux Servers with NVIDIA V100 and NVIDIA A100 (40G) graphics cards with CUDA 11.7.
Appendix E More Ablation Studies
E.1 Computation overhead
Table 18: Training Computational Overhead. We count the average graph size of PubChem and HiPubChem, where HiPubChem adds 9 additional tokens on average. The real preprocessing time and training time are shown below, which are estimated based on 4 A100 40G GPUs. Although HIGHT requires more time to train, the absolute computational overhead of HIGHT is not high.
Graph Size Preprocessing Time Training Time
PubChem 34.39 16min 32sec 8hour 17min 59sec
HiPubChem 43.21 25min 35sec 15hour 36min 23sec
Table 19: Inference Computational Overhead. The inference computational overhead is estimated based on 4 A100 40G GPUs. During the inference, the LLM latency takes up the majority of time. A well-trained LGLM with HIGHT is able to generate more concise and valid answers and thus may take less time during inference.
Property Prediction MolCaption Reagent Prediction Forward Reaction Retrosynthesis
InstructMol 14min 54sec 6hour 22min 27sec 56min 56sec 1hour 34min 28sec 1hour 50min 47sec
HIGHT 15min 12sec 4hour 59min 50sec 50min 29sec 1hour 22min 08sec 1hour 49min 42sec
Table 20: Number of Tunable Parameters during Training. When pretraining the GNN tokenizer, the number of tunable parameters is the number of parameters in GNN encoder; In stage 1, the number of tunable parameters is the number of parameters in the projector; In stage 2, the number of tunable parameters is the number of parameters in the projector and in LoRA.
graph token dimension GNN encoder params in projector params in tokenizer LoRA
InstructMol 300d 1,860,905 1,232,896 3,093,801 159,907,840
HIGHT 300d 1,865,105 3,796,992 5,662,097 159,907,840
E.2 Ablation studies with different setups of the tokenizers
In Table 21, we present more results of the ablation studies with different setups of HIGHT and node-centric tokenizer.
Table 21: More results of chemical reaction tasks with ablation studies. These tasks encompass reagent prediction, forward reaction prediction, and retrosynthesis. $\dagger$: few-shot ICL results from (Fang et al., 2024). *: use task-specific instruction data to finetune.
[…]
Reagent Prediction
Baize (Xu et al., 2023) 0.000 0.051 30.628 0.022 0.018 0.004 0.099
ChatGLM (Zeng et al., 2023) 0.000 0.019 29.169 0.017 0.006 0.002 0.074
InstructMol-G 0.031 0.429 31.447 0.389 0.249 0.220 1.000
+Positional Encoding 0.009 0.423 30.833 0.370 0.231 0.197 0.986
+HiPubChem 0.016 0.473 30.455 0.369 0.237 0.194 0.990
+Large Tokenizer 0.040 0.454 29.163 0.416 0.284 0.248 1.000
InstructMol-GS 0.057 0.439 29.757 0.437 0.314 0.271 0.999
InstructMol+LLama-2-7b-chat 0.016 0.459 29.238 0.359 0.225 0.189 0.988
HIGHT-G 0.050 0.462 28.970 0.441 0.314 0.275 1.000
HIGHT-GS 0.067 0.482 27.167 0.462 0.346 0.303 1.000
HIGHT +LLama-2-7b-chat 0.057 0.495 26.591 0.453 0.333 0.293 1.000
Forward Reaction Prediction
Baize (Xu et al., 2023) 0.000 0.044 41.500 0.004 0.025 0.009 0.097
ChatGLM (Zeng et al., 2023) 0.000 0.183 40.008 0.050 0.100 0.044 0.108
InstructMol-G 0.031 0.853 24.790 0.512 0.362 0.303 0.993
+Positional Encoding 0.0102 0.829 26.622 0.419 0.328 0.268 0.981
+HiPubChem 0.011 0.819 26.010 0.396 0.315 0.264 0.975
+Large Tokenizer 0.040 0.861 24.051 0.544 0.380 0.328 0.996
InstructMol-GS 0.252 0.926 17.773 0.755 0.599 0.543 1.000
InstructMol+LLama-2-7b-chat 0.020 0.841 25.109 0.426 0.339 0.284 0.998
HIGHT-G 0.037 0.869 23.759 0.590 0.394 0.340 0.993
HIGHT-GS 0.293 0.935 16.687 0.774 0.618 0.566 1.000
HIGHT +LLama-2-7b-chat 0.042 0.873 23.854 0.590 0.402 0.344 0.996
Retrosynthesis
Baize (Xu et al., 2023) 0.000 0.095 44.714 0.025 0.050 0.023 0.112
ChatGLM (Zeng et al., 2023) 0.000 0.117 48.365 0.056 0.075 0.043 0.046
InstructMol-G 0.001 0.835 31.359 0.447 0.277 0.241 0.996
+Positional Encoding 0.000 0.793 33.859 0.295 0.218 0.192 0.983
+HiPubChem 0.000 0.755 35.811 0.282 0.218 0.177 0.997
+Large Tokenizer 0.001 0.842 30.613 0.459 0.287 0.263 0.999
InstructMol-GS 0.172 0.911 20.300 0.765 0.615 0.568 1.000
InstructMol+LLama-2-7b-chat 0.000 0.806 32.128 0.292 0.234 0.202 0.985
HIGHT-G 0.008 0.863 28.912 0.564 0.340 0.309 1.000
HIGHT-GS 0.202 0.914 20.194 0.772 0.623 0.577 0.999
HIGHT +LLama-2-7b-chat 0.006 0.865 28.964 0.563 0.338 0.306 0.999
Table 22: Full results of motif hallucinations on MotifHallu with ablation studies.
Method F1 (pos) $\uparrow$ F1 (neg) $\uparrow$ F1 (avg) $\uparrow$
Node-centric Tokenization
InstructMol-G 95.7 9.5 52.6
InstructMol-G+LLama-2-7b-chat 99.6 2.8 51.2
InstructMol-GS 97.1 10.6 53.8
Hierarchical Tokenization
HIGHT-G 85.5 48.2 66.9
HIGHT-G+LLama-2-7b-chat 55.1 65.2 60.2
HIGHT-GS 84.5 42.7 63.6
Ablation variants
InstructMol-G + Positional Encoding 96.4 19.8 58.1
InstructMol-G + HiPubChem 96.6 12.5 54.6
HIGHT-G w/o HiPubChem 96.6 12.5 54.6
HIGHT-GS w/o HiPubChem 98.2 6.5 52.4
Table 23: Results of molecular property prediction tasks (regression) on QM9 with ablation studies. We report the result in MAE. $\dagger$: few-shot in-context learning (ICL) results from (Fang et al., 2024). $\Delta{\epsilon}$ refers to the HOMO-LUMO energy gap.
LLM Based Generalist Models
Alpaca (Dubois et al., 2023) - - - 322.109
Baize (Xu et al., 2023) - - - 261.343
Mol-Instruction (Fang et al., 2024) 0.0210 0.0210 0.0203 0.0210
InstructMol-G 0.0111 0.0133 0.0147 0.0130
+Positional Encodings 0.0300 0.0395 0.0357 0.0350
+HiPubChem 0.0305 4.4019 0.0494 1.1226
HIGHT-G 0.0078 0.0086 0.0095 0.0086